Fermi National Laboratory
 

Operation Summary for September 26, 2011 - September 28, 2011

 

Monday September 26

The day shift began with Operations monitoring Tevatron (TEV) store 9153 (191.4E30) and the Recycler stash (104.5E10), with the Antiproton Source (Pbar) stacking (23.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Booster personnel began conducting Ping and BLM studies at 12:19 PM. Neutron Therapy Facility personnel conducted a Dosimetry run all shift.

During the evening shift, Operations established beam to the Meson FTBF T-992 experiment at 4:30 PM and disabled their beam at 6 PM. At 6:07 PM, the contactors for two Linac Klystron RF stations (KRF2 and 3) opened. KRF3 reset, but KRF2 would not; operators investigated and brought the station on locally. Operations terminated store 9153 at 9:53 PM and put the TEV into shot setup. At 10:35 PM, the Pelletron tripped off; it reset. Operations established store 9154 at 11:14 PM with an average initial luminosity (AIL) of 330E30. Operators reported short down times due to Klystron Vernier trips and MI-8 collimator beam aborts.

Tuesday September 27

The midnight shift began with Operations monitoring store 9154 (245.8E30) and the stash (50.1E10), with Pbar stacking (23.1E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 2:20 AM, the Linac Klystron vernier tripped off; operators reset it. Operators completed a Booster injection tune at 3:07 AM. Operators prepared MTA for access at 7:06 AM. At 7:18 AM, LRF1 and the 750 KeV line suffered a vacuum burst; everything recovered.

During the day shift, at 9:05 AM, a Booster RF vacuum pump (B:IPRF19) tripped off and brought down two Booster RF stations (BRF17 and 18); everything reset and Booster resumed sending beam at 9:44 AM. The MI-30 BPMs tripped off at 9:45 AM for the first of two times in three hours; the BPM system reset. TEV personnel began a Tevatron Electron Lens (TEL2) study at 2:04 PM. Operations established beam to T-922 at 3:39 PM.

During the evening shift, Operations terminated beam to T-922 at 4:32 PM. Operations terminated store 9154 at 4:37 PM, put the TEV into shot setup, and established store 9155 at 5:51 PM with an AIL of 367.1E30.

Wednesday September 28

The midnight shift began with Operations monitoring store 9155 (139.66E30) and the stash (190.7E10), with Pbar stacking (11.3E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, at 12:36 AM, a Booster RF vacuum pump (B:IPRF7) tripped off and brought down BRF7 and 8; everything reset. At 1:30 AM, B:IPRF7 tripped off again; it reset. BRF7 began sparking at 6:02 AM; it reset.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to conduct a study. MiniBooNE personnel reported that they had surpassed 1E21 Protons On Target in antineutrino mode.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/28/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory